A Body of mass m accelarates uniformly from rest to a speed Vo in time To. The work done on the body till any time t is 
  1. 1/2mVo^2(T^2/To^2)
  2. 1/2mVo^2(To/T)
  3. mVo^2(T/To)
  4. mVo^2(T/To)^3
And please explain the steps. 

Grade:12

2 Answers

Saurabh Kumar
9 years ago
Calculate for the power, which is a function of force and velocity,
Since,
Power, P = F .v = (ma)v

From th relations v = u + at = a t (u = 0)
So, P = m a2 t = m (v2/T 2)t

Good Luck..................................
Swagat Bordoloi
9 years ago
Thank you very much...but the answer in the bansal book provided by askiitians has the answer as a0 1/2mv^2t^2/T^2..